# Some Super-approximation Rates of ReLU Neural Networks for Korobov Functions

### v1

@@ 3.1.1 The case $m=2$
where $C_{1}=112d(2d)^{d}$ and $C_{2}=320d^{2}$. Finally, combining $n=\left\lceil 2\log(2WL)\right\rceil$ with (22) and (7) implies
@@
where $C_{3}=(2d)^{3d}$.
@@ 3.1.2 The case $m\geq 3$
$\Pi_{n}^{m}f(\bm{x})$ $=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x})$ (24)
$:=\sum_{|\bm{l}|_{1}\leq n+d-1}f_{\bm{l}}^{m}.$
Given $1\leq p<\infty$, the interpolation error of $\Pi_{n}^{m}f$ has been discussed (see Bungartz and Griebel (2004); Li and Zhang (2025)) and is presented as follows
$\|f-\Pi_{n}^{m}f\|_{L_{p}(\Omega)}\leq C_{m,d}2^{-mn}n^{d},$ (25)
where $C_{m,d}$ is a constant depending only on $m$ and $d$. Furthermore, it is observed in Li and Zhang (2025) that the basis function in (24) could be written as the following form
@@
with $a_{l_{j},i_{j},k},b_{l_{j},i_{j},k}\in\mathbb{R}$ and satisfies $0\leq\rho_{l_{j},i_{j},k}(x_{j})\leq 2^{n+d-1}$. We observe that for $m\geq 3$, the hierarchical Lagrange interpolation of $f\in X^{m}_{p}(\Omega)$ (see Bungartz and Griebel (2004)) on sparse grid can result in at most $2^{m-3}$ possibilities for the shape of the basis functions (see (Bungartz and Griebel, 2004, Section 4.2) and Figure 4.8 therein for instance). We choose the same shape in each support set of the basis function, e.g., the shape of the basis function at the position $\bm{i}=\bm{1}$. Since the $L_{p}$ norm of $\phi_{\bm{l},\bm{i}}^{m}$ is independent of the position of $\bm{i}$, the $L_{p}$-norm of the changed basis functions on $\Omega_{\bm{l},\bm{i}}$ and the error of the sparse grid interpolation remains the same. Henceforth, we will continue to denote the modified basis functions as $\phi_{\bm{l},\bm{i}}^{m}$. And, the product factors could be represented as
@@
for $j=1,2,\ldots,d$ and for $k=1,2\ldots,m-1$. Here, some $c_{k,j}$ values are appropriately set to 00 based on the degree of the basis functions. Since each $\rho_{l_{j},i_{j},k}(x_{j})$ is a piecewise linear function activated by the ReLU function, we can define an affine map as follows
@@
$\sigma\circ\mathcal{B}_{1}\circ g_{1}\circ\mathcal{A}_{2}(\bm{x})=$ $\left(\psi_{l_{1},i_{1},1}(x_{1}),\ldots,\psi_{l_{1},i_{1},m-1}(x_{1}),\ldots,\right.$
$\left.\psi_{l_{d},i_{d},1}(x_{d}),\ldots,\psi_{l_{d},i_{d},m-1}(x_{d}),\bm{x}^{\top}\right)^{\top},$
where $0\leq\psi_{l_{j},i_{j},k}=2^{-n-d+1}\rho_{l_{j},i_{j},k}\leq 1$ for any $j=1,2,\ldots,d$ and $k=1,2,\ldots,m-1$. Here $\mathcal{A}_{2}$ and $g_{1}$ are defined in (14) and (15) respectively, and $\sigma$ represents the ReLU function. By Proposition 4.1 in Lu et al. (2021) with $d=m-1$ and $k=m$, replacing $L$ by $dL$, there exists a function
@@
$|P(\bm{z})-z_{1}z_{2}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}z_{m-1}|\leq 9m(W+1)^{-14mdL}$
for any $\bm{z}=(z_{1},z_{2},\ldots,z_{m-1})\in[0,1]^{m-1}$. Combining this result with Lemma 2 and (16) implies that there exists a function $\Theta_{1}\in\mathcal{NN}(21mdW,28m^{2}dL)$ such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$, the equality
@@
holds. Continuing this process $d-1$ times, we derive a function $\Theta\in\mathcal{NN}(21mdW,28m^{2}d^{2}L)$ such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ we have
@@
$\Big{|}\phi^{0}_{\bm{l}}-\widetilde{v}^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\widetilde{\psi}_{l_{j},i_{j}}(x_{j})\Big{|}$ $\leq C_{m,d}(W+1)^{-14mdL}$ (28)
@@
$|\widetilde{v}^{m}_{\bm{l},\bm{i}}-v^{m}_{\bm{l},\bm{i}}|\leq 2C_{m,d}W^{-2md-2m}(2L)^{-2md-2m}.$ (29)
Then define $\phi^{1}_{\bm{l}}:=2^{md(n+d-1)}\phi^{0}_{\bm{l}}$ for each $\bm{l}$ which satisfies $|\bm{l}|_{1}\leq n+d-1$. Combining (26), (28),(29), Lemma 7, and the inequality $2^{md(n+d-1)}\leq W^{2md}(2L)^{2md}$ implies that
$\Big{|}\phi_{\bm{l}}^{1}-\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x})\Big{|}$ (30)
$\leq$ $2^{md(n+d-1)}\Big{|}\phi_{\bm{l}}^{0}-\widetilde{v}^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\widetilde{\psi}_{l_{j},i_{j}}(x_{j})\Big{|}$
$+2^{md(n+d-1)}\Big{|}\widetilde{v}^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\widetilde{\psi}_{l_{j},i_{j}}(x_{j})-v^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\prod_{k=1}^{m-1}\psi_{l_{j},i_{j},k}(x_{j})\Big{|}$
$\leq$ $C_{m,d}W^{2md}(2L)^{2md}\cdot\Big{(}(W+1)^{-14mdL}$
$+\left.2W^{-2md-2m}(2L)^{-2md-2m}\right.$
@@
By utilizing Lemma 4, (23), and (27), the size of the network function $\phi^{1}$ satisfies
@@
where $C_{7}=70md^{d+1}$, $C_{8}=90m^{2}d^{2}2^{d}$. The remaining process involves estimating the error bounds of neural networks in the $L_{p}$ norm. In this scenario, we can transfer the $L_{p}$ integral of the error from the region $\Omega_{\varepsilon}^{\bm{l}}$ to the entire region $\Omega$ using the same method as in the case where $m=2$. Finally, by combining (25) and (30) and using $n=\left\lceil 2\log(2WL)\right\rceil$, we arrive at
@@ 4 Approximation by ReLU DNN in $W^{1}_{p}$ Norm
The proof of Theorem 2 relies on several concepts and lemmas listed below.
Definition 3.
@@
$g_{1}(x):=\left\{\begin{array}[]{ll}1,&x\in\left[\frac{i}{K}+\frac{1}{4K},\frac{i}{K}+\frac{1}{2K}\right]\\ 0,&x\in\left[\frac{i}{K}+\frac{3}{4K},\frac{i+1}{K}\right]\\ 4K\left(x-\frac{i}{K}\right),&x\in\left[\frac{i}{K},\frac{i}{K}+\frac{1}{4K}\right]\\ -4K\left(x-\frac{4i+3}{4K}\right),&x\in\left[\frac{i}{K}+\frac{1}{2K},\frac{i}{K}+\frac{3}{4K}\right]\end{array},\right.$ (31)
@@
Lemma 8 (Proposition 1 from Yang et al. (2023b)).
@@
where $g_{\bm{k}}$ is determined by $K$ and $d$ as specified in (31).
@@
Lemma 9 (Lemma 7 from Yang et al. (2023b)).
For any $g(\bm{x})\in W^{1}_{\infty}(\Omega)$ and $\phi_{\bm{k}}(\bm{x})$ derived from Lemma 8 with $\bm{k}=\left(k_{1},k_{2},\ldots,k_{d}\right)\in\{1,2\}^{d}$, let $M=\max\left\{\|g\|_{W^{1}_{\infty}\left(\Omega\right)},\left\|\phi_{\bm{k}}\right\|_{W^{1}_{\infty}\left(\Omega\right)}\right\}$, then we have
@@
$\left\|\phi_{\bm{k}}(\bm{x})\cdot g(\bm{x})-\phi_{M}\left(\phi_{\bm{k}}(\bm{x}),g(\bm{x})\right)\right\|_{W^{1}_{\infty}\left(\Omega\right)}$
$=$ $\left\|\phi_{\bm{k}}(\bm{x})\cdot g(\bm{x})-\phi_{M}\left(\phi_{\bm{k}}(\bm{x}),g(\bm{x})\right)\right\|_{W^{1}_{\infty}\left(\Omega_{\bm{k}}\right)}$
where $\phi_{M}$ is from Lemma 5 by choosing $a=M$.
Proposition 1.
@@
4.1 Proof of Theorem 2
In this subsection, we provide the proof of Theorem 2 by using Lemma 5, Lemma 8, Lemma 9, and Proposition 1.
Proof of Theorem 2.
Given $m\geq 2$, by utilizing Lemma 8 and Proposition 1, there exists a sequence of
@@ 4.1 Proof of Theorem 2
$\|\phi_{\bm{k}}(\bm{x})-g_{\bm{k}}(\bm{x})\|_{W^{1}_{\infty}\left(\Omega\right)}\leq 50d^{\frac{5}{2}}(W+1)^{-4dmL},$
for each $\bm{k}\in\{1,2\}^{d}$. Using Lemma 5 with $a=C_{m,d}+50d^{\frac{5}{2}}+1$, there exists a function $\widetilde{\phi}\in\mathcal{NN}(15(W+1),8mL)$ such that
@@
$\leq\Big{\|}\sum_{\bm{k}\in\{1,2\}^{d}}\left(fg_{\bm{k}}-\phi_{\bm{k}}\cdot\psi_{\bm{k}}\right)\Big{\|}_{W^{1}_{p}(\Omega)}$
$+\Big{\|}\sum_{\bm{k}\in\{1,2\}^{d}}\left(\phi_{\bm{k}}\cdot\psi_{\bm{k}}-\widetilde{\phi}(\phi_{\bm{k}},\psi_{\bm{k}})\right)\Big{\|}_{W^{1}_{p}(\Omega)}$
$:=\mathcal{E}_{1}+\mathcal{E}_{2}.$
As for $\mathcal{E}_{1}$, by equation (32), Lemma 9, and the fact
$\|\phi_{\bm{k}}\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ $\leq\|\phi_{\bm{k}}\|_{W^{1}_{\infty}(\Omega)}$
$\leq\left\|\phi_{\bm{k}}(\bm{x})-g_{\bm{k}}\right\|_{W^{1}_{\infty}(\Omega)}+\|g_{\bm{k}}\|_{W^{1}_{\infty}(\Omega)}$
$\leq C_{d}W^{2}L^{2}$
@@
$\leq\sum_{\bm{k}\in\{1,2\}^{d}}C_{p,d}\left(\|f\|_{W^{1}_{p}(\Omega)}\|g_{\bm{k}}-\phi_{\bm{k}}\|_{W^{1}_{\infty}(\Omega)}\right.$
$\quad\left.+\|f-\psi_{\bm{k}}\|_{W^{1}_{p}(\Omega_{\bm{k}})}\|\phi_{\bm{k}}\|_{L_{\infty}(\Omega_{\bm{k}})}\right.$
@@
$\leq\sum_{\bm{k}\in\{1,2\}^{d}}C_{p,d}\left(\|f\|_{W^{1}_{p}(\Omega)}50d^{\frac{5}{2}}(W+1)^{-4dmL}\right.$
$\quad\left.+W^{-2(m-1)}L^{-2(m-1)}(\log(2WL))^{3d}\right)$
@@
As for $\mathcal{E}_{2}$, by equation (32) and Lemma 9, we have
@@
The proof is complete. ∎
In the remaining part of this section, we will prove Proposition 1 below.
4.2 Proof of Proposition 1
@@
for each $\bm{i}=(i_{1},i_{2},\ldots,i_{d})\in\mathcal{I}_{\bm{l}}$ which is defined in (6).

### v2

@@ 3.1.1 The case $m=2$
where $C_{1}=112d(2d)^{d}$ and $C_{2}=320d^{2}$. Finally, combining $n=\left\lceil 2\log(2WL)\right\rceil$ with (3.22) and (2.7) implies
@@
where $C_{3}=(2d)^{3d}$. Substituting $W$ and $L$ with $\frac{W}{C_{1}\log(8W)^{d+1}}$ and $\frac{L}{C_{2}\log(8L)^{d+1}}$, respectively, yields $\phi\in\mathcal{NN}(W,L)$ and the following error bound:
$\|f-\phi\|_{L_{p}(\Omega)}\leq\widetilde{C}_{1}W^{-4}L^{-4}(\log(8W))^{7d+4}(\log(8L))^{7d+4}$
where $\widetilde{C}_{1}=(C_{1}C_{2})^{4}C_{3}$.
@@ 3.1.2 The case $m\geq 3$
$\Pi_{n}^{m}f(\bm{x})$ $=\sum_{|\bm{l}|_{1}\leq n+d-1}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x})=\sum_{|\bm{l}|_{1}\leq n+d-1}f_{\bm{l}}^{m}.$ (3.24)
Given $1\leq p<\infty$, the interpolation error of $\Pi_{n}^{m}f$ has been discussed (see [9, 22]) and is presented as follows
$\|f-\Pi_{n}^{m}f\|_{L_{p}(\Omega)}\leq C_{m,d}2^{-mn}n^{d},$ (3.25)
where $C_{m,d}$ is a constant depending only on $m$ and $d$. Furthermore, it is observed in [22] that the basis function in (3.24) could be written as the following form
@@
with $a_{l_{j},i_{j},k},b_{l_{j},i_{j},k}\in\mathbb{R}$ and satisfies $0\leq\rho_{l_{j},i_{j},k}(x_{j})\leq 2^{n+d-1}$. We observe that for $m\geq 3$, the hierarchical Lagrange interpolation of $f\in X^{m}_{p}(\Omega)$ (see [9]) on sparse grid can result in at most $2^{m-3}$ possibilities for the shape of the basis functions (see [9, Section 4.2] and Figure 4.8 therein for instance). We choose the same shape in each support set of the basis function, e.g., the shape of the basis function at the position $\bm{i}=\bm{1}$. Since the $L_{p}$ norm of $\phi_{\bm{l},\bm{i}}^{m}$ is independent of the position of $\bm{i}$, the $L_{p}$-norm of the changed basis functions on $\Omega_{\bm{l},\bm{i}}$ and the error of the sparse grid interpolation remains the same. Henceforth, we will continue to denote the modified basis functions as $\phi_{\bm{l},\bm{i}}^{m}$. And, the product factors could be represented as
@@
for $j=1,2,\ldots,d$ and for $k=1,2\ldots,m-1$. Here, some $c_{k,j}$ values are appropriately set to 0 based on the degree of the basis functions. Since each $\rho_{l_{j},i_{j},k}(x_{j})$ is a piecewise linear function activated by the ReLU function, we can define an affine map as follows
@@
$\sigma\circ\mathcal{B}_{1}\circ g_{1}\circ\mathcal{A}_{2}(\bm{x})=$ $\left(\psi_{l_{1},i_{1},1}(x_{1}),\ldots,\psi_{l_{1},i_{1},m-1}(x_{1}),\right$.
$\left.\ldots,\psi_{l_{d},i_{d},1}(x_{d}),\ldots,\psi_{l_{d},i_{d},m-1}(x_{d}),\bm{x}^{\top}\right)^{\top},$
where $0\leq\psi_{l_{j},i_{j},k}=2^{-n-d+1}\rho_{l_{j},i_{j},k}\leq 1$ for any $j=1,2,\ldots,d$ and $k=1,2,\ldots,m-1$. Here $\mathcal{A}_{2}$ and $g_{1}$ are defined in (3.14) and (3.15) respectively, and $\sigma$ represents the ReLU function. By Proposition 4.1 in [26] with d=m1d=m-1 and $k=m$, replacing $L$ by $dL$, there exists a function
@@
$|P(\bm{z})-z_{1}z_{2}\cdots z_{m-1}|\leq 9m(W+1)^{-14mdL}$
for any $\bm{z}=(z_{1},z_{2},\ldots,z_{m-1})\in[0,1]^{m-1}$. Combining this result with Lemma 2.2 and (3.16) implies that there exists a function $\Theta_{1}\in\mathcal{NN}(21mdW,28m^{2}dL)$ such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$, the equality
@@
holds. Continuing this process d1d-1 times, we derive a function $\Theta\in\mathcal{NN}(21mdW,28m^{2}d^{2}L)$ such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ we have
@@
$\Big|\phi^{0}_{\bm{l}}-\widetilde{v}^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\widetilde{\psi}_{l_{j},i_{j}}(x_{j})\Big|$ $\leq C_{m,d}(W+1)^{-14mdL}$ (3.28)
@@
$|\widetilde{v}^{m}_{\bm{l},\bm{i}}-v^{m}_{\bm{l},\bm{i}}|\leq 2C_{m,d}W^{-2md-2m}(2L)^{-2md-2m}.$ (3.29)
Then define $\phi^{1}_{\bm{l}}:=2^{md(n+d-1)}\phi^{0}_{\bm{l}}$ for each $\bm{l}$ which satisfies $|\bm{l}|_{1}\leq n+d-1$. Combining (3.26), (3.28), (3.29), Lemma 2.9, and the inequality $2^{md(n+d-1)}\leq C_{m,d}W^{2md}(2L)^{2md}$ implies that
$\Big|\phi_{\bm{l}}^{1}-\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi_{\bm{l},\bm{i}}^{m}(\bm{x})\Big|$ (3.30)
$\leq$ $2^{md(n+d-1)}\left(\Big|\phi_{\bm{l}}^{0}-\widetilde{v}^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\widetilde{\psi}_{l_{j},i_{j}}(x_{j})\Big|+\Big|\widetilde{v}^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\widetilde{\psi}_{l_{j},i_{j}}(x_{j})-v^{m}_{\bm{l},\bm{i}}\prod_{j=1}^{d}\prod_{k=1}^{m-1}\psi_{l_{j},i_{j},k}(x_{j})\Big|\right)$
$\leq$ $C_{m,d}W^{2md}(2L)^{2md}\cdot\Big((W+1)^{-14mdL}+\left.2W^{-2md-2m}(2L)^{-2md-2m}\right$.
@@
By utilizing Lemma 2.4, (3.23), and (3.27), the size of the network function $\phi^{1}$ satisfies
@@
where $C_{7}=70md^{d+1}$, $C_{8}=90m^{2}d^{2}2^{d}$. The remaining process involves estimating the error bounds of neural networks in the $L_{p}$ norm. In this scenario, we can transfer the $L_{p}$ integral of the error from the region $\Omega_{\varepsilon}^{\bm{l}}$ to the entire region $\Omega$ using the same method as in the case where $m=2$. Finally, by combining (3.25) and (3.30) and using $n=\left\lceil 2\log(2WL)\right\rceil$, we arrive at
@@
where $C_{m,d}$ is a constant depending only on $m$ and $d$. Substituting $W$ and $L$ with $\frac{W}{C_{7}\log(8W)^{d+1}}$ and $\frac{L}{C_{8}\log(8L)^{d+1}}$, respectively, yields $\phi\in\mathcal{NN}(W,L)$ and the following error bound:
$\|f-\phi^{1}\|_{L_{p}(\Omega)}\leq C_{m,d}W^{-2m}L^{-2m}(\log(8W))^{2(m+2)(d+1)}(\log(8L))^{2(m+2)(d+1)}$
@@
Remark 3.2.
It is worth noting that when $p=\infty$, the term $\mu((\Omega\backslash\Omega_{\varepsilon}))^{1/p}(\|f_{\bm{l}}\|_{L_{\infty}(\Omega)}+\|\phi_{\bm{l}}\|_{L_{\infty}(\Omega)})$ cannot be rendered sufficiently small simply by reducing $\varepsilon$. Consequently, the approximation error on the entire domain $\Omega$ cannot be directly controlled by the error on $\Omega_{\varepsilon}$ as formulated in (3.22). To address this for $f\in X^{m}_{\infty}(\Omega)$, we can adopt an alternative approach as established in [36, 26] and Theorem 2.1 of [26] to derive the corresponding approximation rates.
@@ 4 Approximation by ReLU DNN in $W^{1}_{p}$ Norm
The proof of Theorem 1.3 relies on several concepts and lemmas listed below.
Definition 4.1.
@@
$g_{1}(x):=\left\{\begin{array}[]{ll}1,&x\in\left[\frac{i}{K}+\frac{1}{4K},\frac{i}{K}+\frac{1}{2K}\right]\\ 0,&x\in\left[\frac{i}{K}+\frac{3}{4K},\frac{i+1}{K}\right]\\ 4K\left(x-\frac{i}{K}\right),&x\in\left[\frac{i}{K},\frac{i}{K}+\frac{1}{4K}\right]\\ -4K\left(x-\frac{4i+3}{4K}\right),&x\in\left[\frac{i}{K}+\frac{1}{2K},\frac{i}{K}+\frac{3}{4K}\right]\end{array},\right$. (4.31)
@@
Lemma 4.3 (Proposition 1 from [38]).
@@
where $g_{\bm{k}}$ is determined by $K$ and $d$ as specified in (4.2).
@@
Lemma 4.4 (Lemma 7 from [38]).
For any $g(\bm{x})\in W^{1}_{\infty}(\Omega)$ and $\phi_{\bm{k}}(\bm{x})$ derived from Lemma 4.3 with $\bm{k}=\left(k_{1},k_{2},\ldots,k_{d}\right)\in\{1,2\}^{d}$, let $M=\max\left\{\|g\|_{W^{1}_{\infty}\left(\Omega\right)},\left\|\phi_{\bm{k}}\right\|_{W^{1}_{\infty}\left(\Omega\right)}\right\}$, then we have
@@
$\left\|\phi_{\bm{k}}(\bm{x})\cdot g(\bm{x})-\phi_{M}\left(\phi_{\bm{k}}(\bm{x}),g(\bm{x})\right)\right\|_{W^{1}_{\infty}\left(\Omega\right)}=\left\|\phi_{\bm{k}}(\bm{x})\cdot g(\bm{x})-\phi_{M}\left(\phi_{\bm{k}}(\bm{x}),g(\bm{x})\right)\right\|_{W^{1}_{\infty}\left(\Omega_{\bm{k}}\right)}$
where $\phi_{M}$ is from Lemma 2.6 by choosing $a=M$.
Proposition 4.5.
@@
4.1 Proof of Theorem 1.3
In this subsection, we provide the proof of Theorem 1.3 by using Lemma 2.6, Lemma 4.3, Lemma 4.4, and Proposition 4.5.
Proof 4.6.
Given $m\geq 2$, by utilizing Lemma 4.3 and Proposition 4.5, there exists a sequence of
@@
[…]
for each $\bm{k}\in\{1,2\}^{d}$. Using Lemma 2.6 with $a=C_{m,d}+50d^{\frac{5}{2}}+1$, there exists a function $\widetilde{\phi}\in\mathcal{NN}(15(W+1),8mL)$ such that
@@
$\leq\Big\|\sum_{\bm{k}\in\{1,2\}^{d}}\left(fg_{\bm{k}}-\phi_{\bm{k}}\cdot\psi_{\bm{k}}\right)\Big\|_{W^{1}_{p}(\Omega)}+\Big\|\sum_{\bm{k}\in\{1,2\}^{d}}\left(\phi_{\bm{k}}\cdot\psi_{\bm{k}}-\widetilde{\phi}(\phi_{\bm{k}},\psi_{\bm{k}})\right)\Big\|_{W^{1}_{p}(\Omega)}$
$=\mathcal{E}_{1}+\mathcal{E}_{2}.$
As for $\mathcal{E}_{1}$, by equation (4.32), Lemma 4.4, and the fact
$\|\phi_{\bm{k}}\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ $\leq\|\phi_{\bm{k}}\|_{W^{1}_{\infty}(\Omega)}\leq\left\|\phi_{\bm{k}}(\bm{x})-g_{\bm{k}}\right\|_{W^{1}_{\infty}(\Omega)}+\|g_{\bm{k}}\|_{W^{1}_{\infty}(\Omega)}\leq C_{d}W^{2}L^{2}$
@@
$\leq\sum_{\bm{k}\in\{1,2\}^{d}}C_{p,d}\left(\|f\|_{W^{1}_{p}(\Omega)}\|g_{\bm{k}}-\phi_{\bm{k}}\|_{W^{1}_{\infty}(\Omega)}+\|f-\psi_{\bm{k}}\|_{W^{1}_{p}(\Omega_{\bm{k}})}\|\phi_{\bm{k}}\|_{L_{\infty}(\Omega_{\bm{k}})}\right$.
@@
$\leq\sum_{\bm{k}\in\{1,2\}^{d}}C_{p,d}\left(\|f\|_{W^{1}_{p}(\Omega)}50d^{\frac{5}{2}}(W+1)^{-4dmL}+W^{-2(m-1)}L^{-2(m-1)}(\log(2WL))^{3d}\right)$
@@
As for $\mathcal{E}_{2}$, by equation (4.32) and Lemma 4.4, we have
@@
Substituting $W$ and $L$ with $\frac{W}{C_{6}(\log 8W)^{d+1}}$ and $\frac{L}{C_{7}(\log 8L)^{d+1}}$, respectively, yields $\phi\in\mathcal{NN}(W,L)$ and the following error bound:
$\|f-\phi\|_{L_{p}(\Omega)}\leq\widetilde{C}_{2}W^{-2(m-1)}L^{-2(m-1)}(\log(8W))^{2(m+1)(d+1)}(\log(8L))^{2(m+1)(d+1)}$
where $\widetilde{C}_{2}$ is a constant depending only on $m,p$ and $d$. The proof is complete.
In the remaining part of this section, we will prove Proposition 4.5 below.
4.2 Proof of Proposition 4.5
@@
for each $\bm{i}=(i_{1},i_{2},\ldots,i_{d})\in\mathcal{I}_{\bm{l}}$ which is defined in (2.6).